# GRACE-MoE: Grouping and Replication with Locality-Aware Routing for Efficient Distributed MoE Inference

### v2

@@ 1 Introduction
Each MoE layer requires two rounds of All-to-All communication, first dispatching tokens to experts and then aggregating results back. Executed repeatedly across layers, this process continuously amplifies communication latency, making communication the primary bottleneck in SMoE inference (He et al., 2022; Gale et al., 2023). In cross-node scenarios, where bandwidth is limited, All-to-All can account for over 70% of the time within a single MoE layer and around 40% of the overall end-to-end inference latency (Li et al., 2023a; Hwang et al., 2023). In parallel, the gating network naturally skews token routing, creating “hot” and “cold” experts. This phenomenon leads to imbalanced computational load, which overloads some GPUs while leaving others underutilized (Lewis et al., 2021; Clark et al., 2022; He et al., 2022). Overloaded GPUs slow down inference, while idle GPUs waste substantial computing resources. Since parallel inference performance is bounded by the slowest device, such imbalance further exacerbates communication tail latency (Go & Mahajan, 2025). More critically, prior work typically addresses only one of these two issues, yet optimizing one often worsens the other. For instance, reducing communication overhead usually aggravates computational load imbalance. This conflict remains unresolved, and most prior work has focused on single-node, multi-GPU settings. Scaling SMoE inference to multi-node environments remains an under-explored challenge, making a joint solution that simultaneously mitigates both issues particularly critical for cross-node scenarios.
@@ 2 Related Work
Expert Placement and Routing. Existing studies on expert placement (e.g., grouping, replication) and routing have attempted to alleviate the bottlenecks in SMoE inference, including FasterMoE (He et al., 2022), FlexMoE (Nie et al., 2023), Prophet (Wang et al., 2023), ExFlow (Yao et al., 2024), Lazarus (Wu et al., 2024), MoETuner (Go & Mahajan, 2025), C2R (Zhang et al., 2025), Speculative MoE (Li et al., 2025), SwiftMoE (Skiadopoulos et al., 2025), EfficientMoE (Zeng et al., 2025), among others. However, these methods generally target either communication overhead or computational load imbalance, but rarely both simultaneously. For example, C2R (Zhang et al., 2025) reduces communication by grouping experts based on collaboration patterns and restricting the routing space, but this exacerbates load imbalance and degrades accuracy. Moreover, most of these works remain confined to single-node, multi-GPU settings, leaving the joint optimization of communication and load balancing in multi-node scenarios largely unexplored.
@@ 6 Conclusion
7 Ethics Statement
This work does not involve human subjects, sensitive data, or other ethical issues. To the best of our knowledge, it does not raise any concerns as outlined in the ICLR Code of Ethics.
8 Reproducibility Statement
We have taken measures to ensure the reproducibility of our results. Section 4 of the main paper provides a detailed description of GRACE-MoE and the design of each module. Section 5.1 further specifies the models, datasets, baselines, evaluation metrics, hardware and software environment used in our experiments, as well as the implementation details of the framework. Additional information, including model architecture details and the corresponding GPU hardware configurations, is provided in Appendix A.4. The information included in the paper and supplementary materials should be sufficient for an independent researcher to reproduce our work. We plan to make the full source code publicly available to facilitate reproducibility.
@@ Appendix A Appendix
A.1 LLM Usage Statement
In preparing this paper, a large language model (LLM) was used solely for grammar checking and language polishing. The design and implementation of all algorithms and experiments, as well as the analysis of results and conclusions, were independently conducted and written by the authors.
A.2 Algorithm for Controlled Non-uniform Grouping
@@
A.3 Algorithm for Topology-Aware Routing with Load Prediction

### v3

@@ 1 Introduction
Each MoE layer requires two rounds of All-to-All communication, first dispatching tokens to experts and then aggregating results back. Executed repeatedly across layers, this process continuously amplifies communication latency, making communication the primary bottleneck in SMoE inference (He et al., 2022; Gale et al., 2023). In cross-node scenarios, where bandwidth is limited, All-to-All can account for over 70% of the time within a single MoE layer and around 40% of the overall end-to-end inference latency (Li et al., 2023a; Hwang et al., 2023). In parallel, the gating network naturally skews token routing, creating “hot” and “cold” experts. This phenomenon leads to imbalanced computational load, which overloads some GPUs while leaving others underutilized (Lewis et al., 2021; Clark et al., 2022; He et al., 2022). Overloaded GPUs slow down inference, while idle GPUs waste substantial computing resources. Since parallel inference performance is bounded by the slowest device, such imbalance further exacerbates communication tail latency (Go and Mahajan, 2025). More critically, prior work typically addresses only one of these two issues, yet optimizing one often worsens the other. For instance, reducing communication overhead usually aggravates computational load imbalance. This conflict remains unresolved, and most prior work has focused on single-node, multi-GPU settings. Scaling SMoE inference to multi-node environments remains an under-explored challenge, making a joint solution that simultaneously mitigates both issues particularly critical for cross-node scenarios.
@@ 2 Related Work
Expert Placement and Routing. Existing studies on expert placement (e.g., grouping, replication) and routing have attempted to alleviate the bottlenecks in SMoE inference, including FasterMoE (He et al., 2022), FlexMoE (Nie et al., 2023), Prophet (Wang et al., 2023), ExFlow (Yao et al., 2024), Lazarus (Wu et al., 2024), MoETuner (Go and Mahajan, 2025), C2R (Zhang et al., 2025), Speculative MoE (Li et al., 2025), SwiftMoE (Skiadopoulos et al., 2025), EfficientMoE (Zeng et al., 2025), among others. However, these methods generally target either communication overhead or computational load imbalance, but rarely both simultaneously. For example, C2R (Zhang et al., 2025) reduces communication by grouping experts based on collaboration patterns and restricting the routing space, but this exacerbates load imbalance and degrades accuracy. Moreover, most of these works remain confined to single-node, multi-GPU settings, leaving the joint optimization of communication and load balancing in multi-node scenarios largely unexplored.
@@ Appendix A Appendix
A.1 Algorithm for Controlled Non-uniform Grouping
@@
A.2 Algorithm for Topology-Aware Routing with Load Prediction